14.34 g of an oxide of lead was heated in a stream of dry hydrogen, 12.42 g of lead was obtained. Relative atomic masses of oxygen and lead are 16 and 207 respectively.
1. How many grams of oxygen were presthe oxide?
2. Calculate the number of moles of oxygen atoms present.
1
Expert's answer
2017-02-01T13:14:11-0500
PbOâ‚‚ + Hâ‚‚ -> Pb + Hâ‚‚O
1Mol 1Mol

n(PbOâ‚‚)=14,34/239=0,06Mol
n(Pb)=12,42/207=0,06Mol
PbOâ‚‚ reacted completely
PbOâ‚‚ -> 2O
239g/mol/14,34g=32g/mol/x
m(O)=1,92g
n(O)=m/n=1,92g/16g/mol=0,12mol